News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Thursday 21 October 1999 11.00am EST

During the last 24 hours Chandra completed observations pf PKS 2155-304 and 4U 1636-53 with ACIS-S and HETG, and of A697 with ACIS-I. The Observatory passed into the radiation belts at 6AM 10/21/99 and will perform post-belt ACIS CTI measurements starting at 2.30PM. A long observation of Abell 370 with ACIS-S will then continue through 10PM 10/22.

Two bias dumps were completed with real-time commanding to help diagnose the ACIS FEP-0 parity error problem. Analysis of the data is underway. In the meantime, observations are continuing with the re-configured ACIS parameter blocks.

Analysis of ACIS CTI data from the last several belt passages has indicated a possible slight increase in the CTI degradation of the front-side illuminated chips. If confirmed, one possible cause is the timing of the safing of the ACIS prior to belt entry. For future loads the mission schedule will ensure that the ACIS is safed 3 hours earlier than for the currently defined belt entry. Work us underway to assess EPHIN data to characterize the belt entry and exit from on-orbit data.

The spacecraft systems continued to operate nominally.





Return to Status Reports